OpenStudy (anonymous):

is represented by the equation 2y + 3x = 2. is constructed such that || . What is the slope of ?

OpenStudy (anonymous):

Seems like there are parts of your question missing. are you trying to the slope of the line paralell to 2y + 3x = 2

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok. you can find the slope of 2y + 3x =2 by writing the equation in slope intercept form. \(y = mx + b\) Do yo know how?

OpenStudy (anonymous):

yeah thank you

OpenStudy (anonymous):

parallel lines have the same slope, so once you have that slope, it will be your answer

OpenStudy (anonymous):

k

OpenStudy (anonymous):

2y=-3x+2\[2y=-3x+2\]

OpenStudy (anonymous):

y=-3/2x+1

OpenStudy (anonymous):

thats the answer right

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the slope is the answer. that's the coefficient of x, so -3/2

OpenStudy (anonymous):

k